ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 142, 2004

Adelman, O. and N. Enriquez Random walks in random environment: What a single trajectory tells

205

Arad, Z., Y. Erez and M. Muzychuk On standard integral table algebras generated by an element of width 2

285

Baumgartner, U. and G. A. Willis Contraction groups and scales of automorphisms of totally disconnected locally compact groups

221

Birman, J. S. and J. A. Moody Obstructions to trivializing a knot 125
Brink, D. On alternating and symmetric groups as Galois groups  47

Enriquez, N. See O. Adelman

Erez, Y. See Z. Arad

Globevnik, J. Analyticity on families of circles  29

Grenié, L. Poles of L-functions, periods and unitary theta. Part I: Odd to even

93

Hoffman, C. A family of nonisomorphic Markov random fields  345

Janson, S., K. Oleszkiewicz and A. Ruciński Upper tails for subgraph counts in random graphs

61

Mann, A. Groups with few class sizes and the centraliser equality subgroup

367

Masser, D. W. Mixing and linear equations over groups in positive characteristic

189

Moody, J. A. See J. S. Birman

Moretó, A. and J. Sangroniz On the number of conjugacy classes of zeros of characters

163

Muzychuk, M. See Z. Arad

Oleszkiewicz, K. See S. Janson

Pakovich, F. On polynomials orthogonal to all powers of a Chebyshev polynomial on a segmant

273

Ruciński, A. See S. Janson

Růžička, P. Lattices of two-sided ideals of locally matricial algebras and the Γ-invariant problem

1

Sangroniz, J. See A. Moretó

Shelah, S. Quite complete real closed fields  261

Smillie, J. and B. Weiss Minimal sets for flows on moduli space  249

Tahzibi, A. Stably ergodic diffeomorphisms which are not partially hyperbolic

315

Weiss, B. See J. Smillie Willis, G. A. See U. Baumgartner